• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Deur

    Log entry time 16:05:10 on November 7,2009

    Entry number 298115

    keyword=Shift summary

    Starting teh shift with 3.16 uA and the optics target in. With optics runs 25303 and 4831.
    08:05 stop run2 25303 & 4831. Start next ones. Optics runs are short so I won't keep a detailed record on them on this summary.

    9:30. Requested Controlled Access for PVDIS DAQ work.
    2:30 Controled access completed. Requested 10 uA of beam.
    2:35. We have 10.8 uA.
    2:40 Raster patern looks weird. We moved to holey carbon and checked raster there. Patern looks fine. We are going back to the carbon. Raster patern still looks weird. Speculations about the optic target to be displaced. (note we check the raster patern with runs from both left and right HRS. Same results are seen).
    3:10 Prepare for Compton tuning. We move the target to empty. Later, we moved back to holey carbon.
    4:00 End of shift. Still doing compton tuning (no beam).